At room temp, NH3 gas at 1 atm and HCl at P atm are allowed to effuse through identical pinholes to the opp. ends og the glass tube 1 m long and unifrom corss section. A white deposit is seen at a distance of 60 cm from HCl end. What is P?

Grade:11

1 Answers

askiitian.expert- chandra sekhar
14 years ago

Hi Vishrant,

P1=1

P2=P

M1=17

M2=36.5

r1/r2=(P1*√M2)/(P2*√M1)

4/6=(1/P)*√(36.5/17)=1.4653/P

P=3/2 * 1.4653

P=2.2 atm